can you help me solve the problem below using the addition metod:
2x+3y=4 and -2x-3y=-4
thank you!!!!!!!!!!!

Answer by rapaljer(4671)

When you add the two equations together:

2x+3y=4
-2x-3y=-4

you get 0=0, which it ALWAYS TRUE. This means that these two equations are actually the SAME LINE. Some textbooks write this as "the set of all values of (x,y) such that 2x +3y = 4." In set notation: { (x,y)| 2x+3y = 4 }.

R*2 at SCC